OpenStudy (anonymous):

if A + B + C = 180 then sin^2A - sin^2b + sin^2 C = ?

OpenStudy (shubhamsrg):

=>(sina + sinb)(sina - sinb) + sin^2 c =>4 sin(a+b /2) cos(a-b /2) sin(a-b /2) cos(a+b /2) + sin^2 c => now a+b = 180 -c we then have 4 sin( 90 -c/2) cos(a-b /2) sin(a-b /2) cos(90- c/2) + sin^2 c =>4cos(a-b /2) sin(a-b /2) cos(c/2) sin(c/2) + sin^2 c now 2 sinx cosx = sin2x =>sin(a-b)(sinc) +sin^2 c =>sinc ( sin(a-b) + sinc) =>sinc (2 sin(a+c-b /2) cos( a-(b+c))/2) =>2sin c (sin(90-b) cos(90 + a) =>-2sinc cosb sina

OpenStudy (anonymous):

@shubhamsrg the answer should be 2sinA cos B sinc

OpenStudy (anonymous):

@satellite73 help

OpenStudy (anonymous):

@shubhamsrg do you know how to use equation editor??

OpenStudy (anonymous):

@satellite73 plzzz help

OpenStudy (anonymous):

= 2 sin(A+B)cos(A-B) + 2sinC cosC =2sinC cos(A-B)+2sinC cosC =2sinC (cos(A-b) + cos C) =2sin C(cos(A-B) - cos(A+B) ) = 2sinC . 2sin A sin B =4 sinA sin B sin C

OpenStudy (anonymous):

@Rohangrr the answer should be 2sinA cos B sinc

OpenStudy (anonymous):

sorry i thought of PI

OpenStudy (anonymous):

@satellite73 plzzz help plzzz

OpenStudy (experimentx):

sinc ( sin(a-b) + sinc) = sinc ( sin(a-b) + sin(180 - (a+b)) = sinc ( sin(a-b) + sin(a+b)) = 2 sin c sin a cos b